# VLM-AD: End-to-End Autonomous Driving through Vision-Language Model Supervision

### v1

@@ Abstract
Human drivers rely on commonsense reasoning to navigate diverse and dynamic real-world scenarios. Existing end-to-end (E2E) autonomous driving (AD) models are typically optimized to mimic driving patterns observed in data, without capturing the underlying reasoning processes. This limitation constrains their ability to handle challenging driving scenarios. To close this gap, we propose VLM-AD, a method that leverages vision-language models (VLMs) as teachers to enhance training by providing additional supervision that incorporates unstructured reasoning information and structured action labels. Such supervision enhances the model’s ability to learn richer feature representations that capture the rationale behind driving patterns. Importantly, our method does not require a VLM during inference, making it practical for real-time deployment. When integrated with state-of-the-art methods, VLM-AD achieves significant improvements in planning accuracy and reduced collision rates on the nuScenes dataset.
@@ 1 Introduction
End-to-end autonomous driving (AD) unifies perception, prediction, and planning into a single framework. This integration aims to coordinate multiple complex tasks, including detection, tracking, mapping, prediction, and planning. Recent approaches [17, 18, 26] have tackled these challenges by using sensor data to generate planned ego trajectories with a single, holistic model. Although these methods have shown promising results, their performance degrades in challenging, long-tail events [7, 5]. On the other hand, human drivers often handle such scenarios effectively, by reasoning through the driving environment and adapting their actions accordingly. This highlights a training gap in current E2E models, which rely solely on the trajectory supervision as sequences of points, lacking the reasoning information necessary for learning rich and robust feature representations to achieve better driving performance.
Manual annotation of reasoning information is often costly, time-consuming, and prone to inconsistent and subjective results, making it difficult to obtain high-quality and scalable annotations. Large foundation models offer an alternative by providing their reasoning capabilities for complex tasks such as driving. Recent methods [39, 4, 27, 55, 51, 61, 40, 54, 13, 20] have directly integrated large foundation models, such as large language models (LLMs) [12, 44, 52] and vision-language models (VLMs) [45, 31, 32, 36], into AD systems to leverage their reasoning capabilities. However, these methods require extensive fine-tuning to translate language-based outputs into precise numerical results, such as planned trajectories or control signals. In addition, these methods rely on large foundation models during inference, which significantly increases both training costs and inference time, making these methods impractical for real-world applications
Given the limitations of manual annotation and the challenges of directly integrating large foundation models into driving systems, we pose the following question: Can large foundation models, such as VLMs, generate reasoning-based text information to enhance autonomous driving models without requiring integration at inference time?
Motivated by this question, we propose VLM-AD, illustrated in Fig. 1, a novel method that leverages VLMs as teachers to automatically generate reasoning-based text annotations. These annotations then serve as supplementary supervisory signals to train end-to-end pipelines, extending beyond standard trajectory labels. Specifically, given a sequence of multi-view images and the future trajectory of the ego vehicle, we project the future trajectory onto the initial front-view image to incorporate critical temporal movement information. We then prompt the VLM model with targeted questions regarding the vehicle’s current status, intended future actions, and reasoning process to generate both freeform and structured responses, thus infusing critical VLM knowledge into the training pipeline.
@@
  • Extensive experiments on the nuScenes dataset validate the effectiveness of our proposed method, showing significant improvements of 14.6% and 33.3% in L2 planning error, and reductions in collision rate by 38.7% and 57.4% for UniAD and VAD, respectively.
@@ 2 Related Work
End-to-End Autonomous Driving. End-to-end autonomous driving systems jointly train all modules toward a unified goal, resulting in reduced information loss throughout the pipeline. Unified frameworks such as ST-P3 [17] and UniAD [18] propose vision-based end-to-end AD systems that unify perception, prediction, and planning. These models achieve state-of-the-art results on the open-loop nuScenes dataset [3]. Following works, such as VAD [26] and VADv2 [6], introduce a vectorized encoding approach for efficient scene representation and extend to closed-loop simulation on CARLA [14]. Recent methods like Ego-MLP [62], BEV-Planner [35], and PARA-Drive [58] have been developed to explore ego-status and novel design spaces within modular stacks to further enhance driving performance. While E2E driving models show promising results in the development of E2E driving methods, they are primarily optimized to mimic driving patterns in the data, without capturing the underlying reasoning processes. This limitation is largely due to the lack of reasoning information in existing datasets. Consequently, these methods are unable to acquire deeper reasoning knowledge, which could limit their performance in challenging scenarios.
VLMs address this gap by integrating language and vision for multimodal reasoning, supporting tasks like scene understanding [42, 21, 10, 49] and data generation [24, 56, 64]. VLMs have also been used for unified navigation and planning [29, 51, 53, 15] as well as end-to-end autonomous driving [27, 55, 61, 40]. However, existing VLM-based methods often require extensive domain-specific fine-tuning, which significantly increases computational cost and inference latency. Closely related to our method in end-to-end autonomous driving, VLP [40] transforms ground-truth trajectory and bounding box labels into text features for contrastive learning, but it does not introduce information beyond existing supervision labels. In contrast, our method leverages VLMs to provide additional reasoning information to further enhance driving performance.
Multi-Task Learning. Multi-task learning (MTL) jointly performs several related tasks using a shared representation through separate branches or heads. This approach leverages shared domain knowledge, enhancing feature robustness and generalization, making it well-suited for end-to-end autonomous driving. In AD systems, auxiliary tasks such as semantic segmentation [9, 23, 19, 33, 60], depth estimation [33, 60], HD mapping and BEV segmentation [8, 63, 47, 25, 48] are commonly adopted to extract meaningful perception representations for subsequent objects. Beyond vision tasks, other methods [22, 59] predict additional traffic light states or control signals, to improve driving performance. Inspired by the success of multi-task learning, we design novel auxiliary tasks to encourage the model to learn richer feature representations through high-quality reasoning annotations from VLMs, ultimately leading to more reliable planning performance.
3 Method
Fig. 2 presents an overview of our proposed VLM-AD framework, which consists of two main components. The first is the annotation branch, where we leverage the VLM to generate additional information, creating a supplementary dataset that serves as supervision. The second component is our designed auxiliary heads, which aim to align with this additional supervision and can be effectively integrated into any E2E model following the planning module.
@@ 3.1 VLM Text Annotation
Fig. 3 illustrates the annotation process, where we utilize a VLM as the teacher to enrich the dataset with additional information, leveraging its reasoning capabilities from visual inputs to deepen an E2E model’s understanding of driving behaviors. The annotation process can be defined as:
@@
In our work, we employ GPT-4o [2], a high-performance VLM trained on internet-scale data, to automatically annotate our dataset. GPT-4o can interpret the scenario, generate suitable reasoning-based responses, and accurately identify the actions of the ego vehicle in complex scenarios.
@@
Figure 3: Example annotations from GPT-4o. Given consecutive frames from the front-view camera of the ego vehicle, we project the future trajectory of the vehicle onto the first frame. Using our prompts, we obtain three freeform text responses on the current ego status, anticipated actions, and reasoning, along with three structured action responses.
@@
To maximize the reasoning capabilities of the VLM, we provide detailed context descriptions as preliminary instructions before posing specific questions. Specifically, the context and questions are defined as follows:
  • $C_{1}$: This is the front-view image of the ego vehicle. The red line indicates the future trajectory, no line suggests stopping or slowing down. When explaining the reasoning, please focus on the camera image and surrounding context rather than referencing the plotted trajectory.
@@
The complete input prompt is defined as $\mathcal{P}_{1}=[C_{1},Q_{1}]$, where $Q_{1}$ represents the set of questions, $Q_{1}=\{Q_{1-1},Q_{1-2},Q_{1-3}\}$. These open-ended questions yield free-form text annotations describing the ego vehicle’s current status, intended future actions, and the reasoning underlying the VLM’s knowledge.
@@
The complete input prompt is defined as $\mathcal{P}_{2}=[C_{2},Q_{2}]$, where $Q_{2}$ represents the set of structured action questions, $Q_{2}=\{Q_{2-1},Q_{2-2},Q_{2-3}\}$. In this way, we can obtain three specific actions from the VLM. Compared to freeform text annotations, one major benefit of structured annotations is that they can be used to supervise an E2E driving model to predict human-interpretable actions, as demonstrated in the experimental results of Sec. 4.
@@ 3.2 Auxiliary Heads
$y_{1}&=\text{CLIP}(\mathcal{A}_{1})\\ y_{2}&=\text{One-Hot}(\mathcal{A}_{2})\end{aligned},$ (2)
@@
Text Feature Alignment. Using the three text features $y_{1}=\{y_{c},y_{f},y_{r}\}$ as supervision, we develop a feature alignment head that takes the ego feature $f_{\text{ego}}$ as input. This setup resembles knowledge distillation [16], where the feature alignment head learns to align with the text features provided by the teacher VLM.
@@
$q^{\prime}_{1}&=\text{MHCA}(q,k,v)\\ q&=q_{1},k=v=f_{\text{ego}}\\ \hat{f}_{1}&=\text{MLP}(q^{\prime}_{1}\oplus f_{\text{ego}})\\ \end{aligned},$ (3)
@@
Inspired by the knowledge distillation approach in DINO [1] that controls the smoothness and sharpness of feature vectors, we adopt a similar strategy to normalize the text and output features with different temperature parameters, producing feature distributions rather than raw feature values as follows:
[…]
@@
Structured Action Classification. We obtain the structured action labels $y_{2}=\{y_{\text{control}},y_{\text{turn}},y_{\text{lane}}\}$ from the VLM using question $Q_{2}$. We then construct another action classification head that takes the ego feature $f_{\text{ego}}$ as input. Similar to the previous feature alignment stage, we initialize three learnable action queries, $q_{\text{control}}$, $q_{\text{turn}}$, and $q_{\text{lane}}$, which interact with $f_{\text{ego}}$ through three MHCA blocks. In this setup, each action query serves as the attention query $q$, while the ego feature acts as both the key $k$ and value $v$, producing updated action queries. We then concatenate these updated queries with the ego feature to create the feature representation for the action classification head, passing it through an MLP layer followed by a Softmax function to generate the action predictions. This process is formulated as follows:
$q^{\prime}_{2}&=\text{MHCA}(q,k,v)\\ q&=q_{2},k=v=f_{\text{ego}}\\ \hat{f}_{2}&=\text{Softmax}\left(\text{MLP}(q^{\prime}_{2}\oplus f_{\text{ego}})\right)\end{aligned},$ (5)
@@ 3.3 Auxiliary Loss
We define two parallel auxiliary tasks following the planning module to enable the model to distill knowledge from the VLM, and the overall training loss is defined as a weighted sum of two components:
@@
where each component corresponds to a distinct auxiliary text head, providing supervision in the targeted area:
$\mathcal{L}_{\text{align}}=&-P(y_{c})\log(P(\hat{f}_{c}))-P(y_{f})\log(P(\hat{f}_{f}))\\ &-P(y_{r})\log(P(\hat{f}_{r}))\\ \mathcal{L}_{\text{action}}=&-\sum_{i=1}^{N_{\text{control}}}y_{\text{control}}^{i}\log(\hat{f}_{\text{control}}^{i})\\ &-\sum_{i=1}^{N_{\text{turn}}}y_{\text{turn}}^{i}\log(\hat{f}_{\text{turn}}^{i})\\ &-\sum_{i=1}^{N_{\text{lane}}}y_{\text{lane}}^{i}\log(\hat{f}_{\text{lane}}^{i})\end{aligned}.$ (7)
@@
Table 1: Planning results of our VLM-AD method and baselines. The best is in bold and the second best is underlined. VLM-AD consistently outperforms the baselines, with the reasoning-focused $Q_{1}$ contributing the most significant improvements.

### v2

@@ Abstract
Human drivers rely on commonsense reasoning to navigate diverse and dynamic real-world scenarios. Existing end-to-end (E2E) autonomous driving (AD) models are typically optimized to mimic driving patterns observed in data, without capturing the underlying reasoning processes. This limitation constrains their ability to handle challenging driving scenarios. To close this gap, we propose VLM-AD, a method that leverages vision-language models (VLMs) as teachers to enhance training by providing additional supervision that incorporates unstructured reasoning information and structured action labels. Such supervision enhances the model’s ability to learn richer feature representations that capture the rationale behind driving patterns. Importantly, our method does not require a VLM during inference, making it practical for real-time deployment. When integrated with state-of-the-art methods, VLM-AD achieves significant improvements in planning accuracy and reduced collision rates on the nuScenes dataset. It further improves route completion and driving scores under closed-loop evaluation, demonstrating its effectiveness in long-horizon, interactive driving scenarios and its potential for safe and reliable real-world deployment.
Keywords: End-to-End Autonomous Driving, Vision-Language Model, Multimodal Large Language Model, Foundation Models for Driving
@@ 1 Introduction
End-to-end autonomous driving (AD) unifies perception, prediction, and planning into a single framework. This integration aims to coordinate multiple complex tasks, including detection, tracking, mapping, prediction, and planning. Recent approaches [1, 2, 3] have tackled these challenges by using sensor data to generate planned ego trajectories with a single, holistic model. Although these methods have shown promising results, their performance degrades in challenging, long-tail events [4, 5]. On the other hand, human drivers often handle such scenarios effectively by reasoning through the driving environment and adapting their actions accordingly. This highlights a training gap in current E2E models, which rely solely on the trajectory supervision as sequences of points, lacking the reasoning information necessary for learning rich and robust feature representations to achieve better driving performance.
Manual annotation of reasoning information is often costly, time-consuming, and prone to inconsistent and subjective results, making it difficult to obtain high-quality and scalable annotations. Large foundation models offer an alternative by providing their reasoning capabilities for complex tasks such as driving. Recent methods [6, 7, 8, 9, 10, 11, 12, 13, 14, 15] have directly integrated large foundation models, such as large language models (LLMs) [16, 17, 18] and vision-language models (VLMs) [19, 20, 21, 22], into AD systems to leverage their reasoning capabilities. However, these
@@
methods require extensive fine-tuning to translate language-based outputs into precise numerical results, such as planned trajectories or control signals. In addition, these methods rely on large foundation models during inference, which significantly increases both training costs and inference time, making these methods impractical for real-world applications. Given the limitations of manual annotation and the challenges of directly integrating large foundation models into driving systems, we pose the following question: Can large foundation models, such as VLMs, generate reasoning-based text information to enhance autonomous driving models without requiring integration at inference time?
Motivated by this question, we propose VLM-AD, illustrated in Fig. 1, a novel method that leverages VLMs as teachers to automatically generate reasoning-based text annotations. These annotations then serve as supplementary supervisory signals to train end-to-end pipelines, extending beyond standard trajectory labels. Specifically, given a sequence of multi-view images and the future trajectory of the ego vehicle, we project the future trajectory onto the initial front-view image to incorporate critical temporal movement information. We then prompt the VLM model with questions regarding the vehicle’s current status, intended future actions, and reasoning process to generate both freeform and structured responses, thus infusing critical VLM knowledge into the training pipeline.
@@
  • Extensive experiments demonstrate significant improvements in L2 planning accuracy and collision rate over UniAD, VAD, and SparseDrive on the nuScenes dataset, as well as higher route completion and driving scores on the closed-loop CARLA Town05 benchmark.
@@ 2 Related Work
End-to-End Autonomous Driving. End-to-end autonomous driving systems jointly train all modules toward a unified goal, resulting in reduced information loss throughout the pipeline. Unified frameworks such as ST-P3 [1] and UniAD [2] propose vision-based end-to-end AD systems that unify perception, prediction, and planning. These models achieve state-of-the-art results on the open-loop nuScenes dataset [23]. Following works, such as VAD [3] and VADv2 [24], introduce a vectorized encoding approach for efficient scene representation and extend to closed-loop simulation on CARLA [25]. Recent methods like Ego-MLP [26], BEV-Planner [27], and PARA-Drive [28] have been developed to explore ego-status and novel design spaces within modular stacks to further enhance driving performance. In parallel, methods like GenAD [29], GraphAD [30], SparseAD [31], and SparseDrive [32] continue to push the boundaries of end-to-end driving systems through improved representations and task formulations. While E2E driving models show promising results in the development of E2E driving methods, they are primarily optimized to mimic driving patterns in the data, without capturing the underlying reasoning processes. This limitation is largely due to the lack of reasoning information in existing datasets. Consequently, these methods are unable to acquire deeper reasoning knowledge, which could limit their performance in challenging scenarios.
VLMs address this gap by integrating language and vision for multimodal reasoning, supporting tasks like scene understanding [34, 35, 36, 37] and data generation [38, 39, 40]. VLMs have also been used for unified navigation and planning [41, 10, 42, 43] as well as end-to-end autonomous driving [8, 9, 11, 12, 44]. However, existing VLM-based methods often require extensive domain-specific fine-tuning, which significantly increases computational cost and inference latency. Closely related to our method in end-to-end autonomous driving, VLP [12] transforms ground-truth trajectory and bounding box labels into text features for contrastive learning, but it does not introduce information beyond existing supervision labels. In contrast, our method leverages VLMs to provide additional reasoning information to further enhance driving performance.
3 Method
@@
Fig. 2 presents an overview of our proposed VLM-AD framework, which consists of two main components. The first is the annotation component, where a VLM generates rich auxiliary information, forming a supplementary dataset for supervision. The second component is our designed auxiliary heads designed to align with this supervision, which can be seamlessly integrated into any E2E model following the planning module.
@@ 3.1 VLM Text Annotation
We utilize a VLM as the teacher to enrich the dataset with additional information, leveraging its reasoning capabilities from visual inputs to deepen an E2E model’s understanding of driving behaviors. The annotation process can be defined as:
@@
In our work, we employ GPT-4o [45], a high-performance VLM trained on internet-scale data, to automatically annotate our dataset. GPT-4o is used to interpret the scenario, generate suitable reasoning-based responses, and identify the actions of the ego vehicle in complex scenarios.
@@
To maximize the reasoning capabilities of the VLM, we provide detailed context descriptions as preliminary instructions before posing specific questions, defined as follows:
  • $C_{1}$: This is the front-view image of the ego vehicle. The red line indicates the future trajectory, no line suggests stopping or slowing down. When explaining the reasoning, please focus on the camera image and the surrounding context rather than referencing the plotted trajectory.
@@
The full input prompt is defined as $\mathcal{P}_{1}=[C_{1},Q_{1}]$, where $Q_{1}=\{Q_{1-1},Q_{1-2},Q_{1-3}\}$ represents the set of questions. These open-ended questions yield free-form text annotations describing the ego vehicle’s current status, intended future actions, and the reasoning underlying the VLM’s knowledge.
@@
The complete input prompt is defined as $\mathcal{P}_{2}=[C_{2},Q_{2}]$, where $Q_{2}=\{Q_{2-1},Q_{2-2},Q_{2-3}\}$ represents the set of structured action questions. In this way, we can obtain three specific actions from the VLM. Compared to freeform text annotations, one major benefit of structured annotations is that they can be used to supervise an E2E driving model to predict human-interpretable actions, as demonstrated in the experimental results of Sec. 4.
@@ 3.2 Auxiliary Heads
$y_{1}=\text{CLIP}(\mathcal{A}_{1}),\quad y_{2}=\text{One-Hot}(\mathcal{A}_{2}),$ (2)
@@
Text Feature Alignment. Using the three text features $y_{1}=\{y_{c},y_{f},y_{r}\}$ as supervision, we develop a feature alignment head that takes the ego feature $f_{\text{ego}}$ as input. This setup resembles knowledge distillation, where the alignment head is trained to match the text features provided by the VLM.
@@
$q^{\prime}_{1}=\text{MHCA}(q,k,v),\quad q=q_{1},k=v=f_{\text{ego}},\quad\hat{f}_{1}=\text{MLP}(q^{\prime}_{1}\oplus f_{\text{ego}}),$ (3)
@@
Inspired by the knowledge distillation approach DINO [49] that controls the smoothness and sharpness of features, we adopt a similar strategy to normalize the text and output features with different temperature parameters, producing feature distributions rather than raw feature values as follows:
$P(y_{1})=\frac{\exp(y_{1}/\tau_{t})}{\sum_{k=1}^{C}\exp(y_{1}^{(k)}/\tau_{t})},\quad P(\hat{f}_{1})=\frac{\exp(\hat{f}_{1}/\tau_{s})}{\sum_{k=1}^{C}\exp(\hat{f}_{1}^{(k)}/\tau_{s})},$ (4)
@@
Structured Action Classification. We obtain the structured action labels $y_{2}=\{y_{\text{control}},y_{\text{turn}},y_{\text{lane}}\}$ from the VLM using question $Q_{2}$. We then construct another action classification head that takes the ego feature $f_{\text{ego}}$ as input. Similar to the previous feature alignment stage, we initialize three learnable action queries, $q_{\text{control}}$, $q_{\text{turn}}$, and $q_{\text{lane}}$, which interact with $f_{\text{ego}}$ through three MHCA blocks. In this setup, each action query serves as the attention query $q$, while the ego feature acts as the key $k$ and value $v$, producing updated action queries. We then concatenate updated queries with the ego feature to create the representation for the action classification head, passing it through an MLP layer followed by a Softmax to generate the action predictions. This process is formulated as:
$q^{\prime}_{2}=\text{MHCA}(q,k,v),\quad q=q_{2},k=v=f_{\text{ego}},\quad\hat{f}_{2}=\text{Softmax}\left(\text{MLP}(q^{\prime}_{2}\oplus f_{\text{ego}})\right),$ (5)
@@ 3.3 Auxiliary Loss
We define two parallel auxiliary tasks following the planning module to enable the model to distill knowledge from the VLM, and the overall training loss is a weighted sum of two components:
@@
where each component corresponds to a distinct auxiliary text head to provide supervision:
$\begin{aligned} \mathcal{L}_{\text{align}}=&-P(y_{c})\log(P(\hat{f}_{c}))-P(y_{f})\log(P(\hat{f}_{f}))-P(y_{r})\log(P(\hat{f}_{r}))\\ \mathcal{L}_{\text{action}}=&-\sum_{i=1}^{N_{\text{control}}}y_{\text{control}}^{i}\log(\hat{f}_{\text{control}}^{i})-\sum_{i=1}^{N_{\text{turn}}}y_{\text{turn}}^{i}\log(\hat{f}_{\text{turn}}^{i})-\sum_{i=1}^{N_{\text{lane}}}y_{\text{lane}}^{i}\log(\hat{f}_{\text{lane}}^{i})\end{aligned}.$ (7)
@@
4 Experiments
4.1 Open-Loop Settings
Baselines. Our proposed method is a general framework compatible with various end-to-end autonomous driving methods. We validate its effectiveness by applying it to three widely recognized open-source methods, UniAD [2], VAD [3], and SparseDrive [32]. Additionally, we compare it with VLP [12], which projects ego vehicle ground-truth labels into text feature space via CLIP [19] for contrastive learning.
Metrics. We focus on the planning task and use standard metrics such as L2 displacement error and collision rate to evaluate performance for open-loop evaluation.
4.2 Closed-Loop Settings
We further evaluate our proposed method in a closed-loop setting. Following the protocol established by VAD [3], we conduct simulations using the CARLA [25] simulator and the Town05 [50] benchmark. To ensure a fair comparison, we integrate our method into the VAD framework and adopt the same evaluation definitions for closed-loop evaluation.
Metrics. We use the standard Route Completion (RC) and Driving Score (DS) metrics for closed-loop performance evaluation.
4.3 Implementation Details
We use official codes of the UniAD, VAD, and SparseDrive, adhering to the hyperparameters specified in their official implementations. For our proposed VLM-AD, we define two auxiliary task heads, each containing an MHCA block with 8 heads and 3 cross-attention layers, and we set 3 text queries for each of $Q_{1}$ and $Q_{2}$. During training, we set the temperature parameters $\tau_{s}=0.1$ and $\tau_{t}=0.04$ to control the sharpness of the features, and we set $\lambda_{1}=1$ and $\lambda_{2}=0.1$ to balance $\mathcal{L}_{align}$ and $\mathcal{L}_{action}$. All models are trained on 8 NVIDIA H100 GPUs using the PyTorch framework [51]. Complete implementation details, annotation quality analysis, and additional experiments, including visualizations, are provided in the appendix.
4.4 Main Results
Table 1: Planning results of our VLM-AD method and baselines. The best results are shown in bold and the second best are underlined. VLM-AD consistently outperforms the baselines, with the reasoning-focused $Q_{1}$ contributing the most significant improvements.
@@ 4.4 Main Results
17 SparseDrive-B [32] 0.29 0.55 0.91 0.58 0.01 0.02 0.13 0.06 Official
18 VLM-AD 0.25 0.50 0.86 0.54 0.01 0.02 0.12 0.05 -
19 VLM-AD 0.27 0.52 0.87 0.55 0.01 0.02 0.13 0.05 -
20 VLM-AD 0.24 0.49 0.84 0.52 0.01 0.02 0.12 0.05 -
Open-Loop Planning Results. Tab. 1 presents the results of applying our proposed VLM-AD to three baselines, UniAD, VAD, and SparseDrive, as well as comparisons with VLP. Comparing methods ID 0 and 1, we achieve nearly identical planning results using the authors’ officially trained checkpoints. For methods IDs 6 and 7, and IDs 12 and 13, we observe some discrepancies between our reproduced results and the reported values, which we attribute to a correction in image configuration in the official codebase [^1] https://github.com/hustvl/VAD/issues/9. From the first section of the table, we observe that VLM-AD significantly outperforms UniAD on both average L2 planning error and average collision rate by introducing $Q_{1}$ and $Q_{2}$. It also surpasses a state-of-the-art baseline VLP in both metrics. Regarding VAD, our VLM-AD consistently outperforms both VAD-Base and VAD-Tiny, particularly on the L2 planning error metric, and it achieves superior performance compared to the VLP in VAD-Base. For SparseDrive-B, we observe that integrating our VLM-AD leads to improvements in L2 planning error over the baseline model, while maintaining comparable or slightly better collision rates. This is largely due to the collision-aware rescore module in the baseline, which already effectively reduces collision rates. Importantly, the integration of VLM-AD does not compromise this strength, further demonstrating the robustness of our approach. These results demonstrate the effectiveness and advantages of our VLM-AD approach. Additionally, $Q_{1}$ yields better results than $Q_{2}$, verifying the value of supervising the driving model through rich reasoning information.
Table 2: Closed-loop evaluation results of our VLM-AD method and baselines. The best is in bold and the second best is underlined. : LiDAR-based method.
Method Town05 Short Town05 Long
DS $\uparrow$ RC $\uparrow$ DS $\uparrow$ RC $\uparrow$
CILRS [52] 7.47 13.40 3.68 7.19
LBC [53] 30.97 55.01 7.05 32.09
Transfuser [50] 54.52 78.41 33.15 56.36
ST-P3 [1] 55.14 86.74 11.45 83.15
VAD-Base [3] 64.29 87.26 30.31 75.20
+VLM-AD 67.78 88.56 35.25 84.14
Closed-Loop Planning Results. Tab. 2 presents the closed-loop results on the Town05 benchmark for both short and long routes. Our proposed method, VLM-AD, achieves the best performance across all metrics, including Driving Score (DS) and Route Completion (RC). It surpasses VAD-Base and ST-P3, two strong vision-based baselines, with noticeable gains particularly in DS on the long route. These results demonstrate the effectiveness and robustness of VLM-AD in realistic closed-loop driving scenarios, highlighting its ability to reason and act reliably under long-horizon tasks.
4.5 Ablation Study
Table 3: Ablation study on the contributions of $Q_{1-1}$, $Q_{1-2}$ and $Q_{1-3}$ to the UniAD baseline. The best is in bold and the second best is underlined.